# Multitask methods for predicting molecular properties from heterogeneous data

### v1

@@ Abstract
Data generation remains a bottleneck in training surrogate models to predict molecular properties. We demonstrate that multitask Gaussian process regression overcomes this limitation by leveraging both expensive and cheap data sources. In particular, we consider training sets constructed from coupled-cluster (CC) and density function theory (DFT) data. We report that multitask surrogates can predict at CC level accuracy with a reduction to data generation cost by over an order of magnitude. Of note, our approach allows the training set to include DFT data generated by a heterogeneous mix of exchange-correlation functionals without imposing any artificial hierarchy on functional accuracy. More generally, the multitask framework can accommodate a wider range of training set structures—including full disparity between the different levels of fidelity—than existing kernel approaches based on $\Delta$-learning, though we show that the accuracy of the two approaches can be similar. Consequently, multitask regression can be a tool for reducing data generation costs even further by opportunistically exploiting existing data sources.
@@ I Introduction
Predicting molecular properties based on first-principle approaches drives innovation across the physical sciences. However, exact prediction is infeasible due to the exponential scaling of the Schrödinger equation—the underlying quantum-mechanical description of materials systems. Decades of research have yielded numerous first-principle models to describe electronic structures approximately, each representing a compromise between computational cost and prediction accuracy. Coupled cluster (CC) methods are typical examples of high fidelity models which offer accuracy in exchange for considerable computational cost. Specifically, CCSD(T)—the “gold standard” of quantum-chemical prediction—suffers from a steep $O(N^{7})$ scaling where $N$ is the number of electrons in the target system. Harding et al. (2008) In contrast, density-functional theory (DFT) methods are popular baseline models since they provide adequate accuracy for most systems while enjoying substantially lower cost (scaling only as $O(N^{3})$). Ying, Yu, and Ying (2019); Perdew and Schmidt (2001); Goerigkab and Grimme (2011); Teale et al. (2022)
@@
Traditional workflows force users to settle on a single model which is cheap enough to result in manageable cost but still sufficiently high fidelity to give meaningful predictions. However, the advent of modern data-driven modeling has introduced new tools which make it possible to decouple expensive data generation from prediction. Statistical surrogates are trained with predictions from first-principle models and, once trained, can be substituted into workflows, making the prediction step considerably cheaper. In the context of building potential energy surfaces and interatomic potentials, these data-driven approaches have become the de facto standard. Smith et al. (2019); Dral et al. (2020); Goodlett, Turney, and Schaefer (2023); Zaverkin et al. (2023) Simultaneously, large-scale data generation is facilitated by sophisticated workflow engines for high-throughput calculations Curtarolo et al. (2012); Jain et al. (2011); Huber et al. (2020) and robust DFT methods inspired by mathematical research. Teale et al. (2022); Cancès et al. (2022); Herbst and Levitt (2020, 2022); Cancès et al. (2023) The increasing amount of openly available data Ruddigkeit et al. (2012); Ramakrishnan et al. (2014); Chanussot* et al. (2021); Smith, Isayev, and Roitberg (2017) could in principle be readily employed, but quality is generally not homogeneous as each dataset’s author has made different modeling choices. Consequently, most surrogate models are single fidelity—all training data is generated using the same quantum-chemical method and the same numerical approach. Thus, heterogeneity has been an obstacle to opportunistically exploiting the already available data. With this work, we contribute to the promising development of multifidelity surrogates which do not require a single choice between accurate (but expensive) and fast (but crude) methods, but instead combine the advantages of both. Specifically, we consider multitask methods which construct a relationship between multiple regression problems without requiring precise information about the respective accuracy of each source. Bonilla, Chai, and Williams (2008); Leen, Peltonen, and Kaski (2012)
Gaussian process (GP) regression is a popular tool for constructing surrogates informed by heterogeneous data. Several works Pilania, Gubernatis, and Lookman (2017); Batra et al. (2019) have explored the use of multifidelity GP models Kennedy and O’Hagan (2000) trained using two fidelities, namely a high and a low fidelity DFT approach. Alternatively, $\Delta$-learning may be employed to model the difference between two levels of theory, then by adding that difference to low level predictions of the target to produce high level predictions. Ramakrishnan et al. (2015); Dral, Zubatiuk, and Xue (2023); Batra et al. (2019) Though GP regression and the related technique of kernel ridge regression are common choices for fitting such models, in principle, any method of fitting a surrogate may be employed. Batra et al. (2019); Dral, Zubatiuk, and Xue (2023). $\Delta$-learning can be extended to leverage more than two training data sets if multiple difference models are learned and summed. Dral et al. (2020); Dral, Zubatiuk, and Xue (2023) However, between each pair of training sets, data points must be in some way aligned to allow for learning a difference model. For example, in a model of the difference between DFT and CCSD(T), DFT predictions may only contribute to the training set if a corresponding CCSD(T) prediction is also available. Further, when either $\Delta$-learning or multifidelity GP regression learn from more than two heterogeneous training sets, these sets must be assigned some order—for example, following a hierarchy in accuracy. Recent work Vinod, Kleinekathöfer, and Zaspel (2023) has proposed an optimized linear combination of submodels as an alternative, but all coefficients in this linear combination must be re-optimized whenever any submodel is altered, and opportunistic data set construction is not fully explored.
Progress towards multifidelity approaches has also been made in the context of neural network surrogates. For example, transfer learning methods train an initial neural network on a larger source dataset, then re-optimise part of the parameters using a target dataset. Towell and Shavlik (1994); Fu (1989); Smith et al. (2019) For the prediction of atomistic properties, the source dataset might contain millions of DFT calculations while the target might be a smaller set of CCSD(T) predictions. In prior studies, it has been observed that transfer learning can produce higher accuracy than training a network on only one of these datasets. Dral et al. (2020); Dral, Zubatiuk, and Xue (2023) While the framework we discuss in this work is not restricted to GP regression strategies, we will focus exclusively on GP methods. In comparison to neural network approaches, where a rigorous mathematical explanation for the low generalization error of well-trained models is still lacking Bartlett, Montanari, and Rakhlin (2021); Lotfi et al. (2022), GP approaches are more amenable to mathematical analysis. Rasmussen and Williams (2006) This comparative transparency simplifies interpretation of the surrogate’s predictive process in the diverse training data configurations we analyze here. Additionally, GP methods have the advantage of yielding accurate predictions even for relatively small training set sizes. While the training cost of GP regression may become large as the size of the training set increases, many strategies for efficient scaling have been developed and can be routinely employed Quinonero-Candela and Rasmussen (2005); Liu et al. (2018); Wilson, Dann, and Nickisch (2015); Cole, Christianson, and Gramacy (2021).
This work explores a GP regression approach which can accommodate an arbitrary number of quantum chemical datasets without an assumption of hierarchy. In particular, we employ the multitask framework Leen, Peltonen, and Kaski (2012), which has been shown to easily incorporate heterogeneous data sources. The notion of heterogeneity in fact encompasses many possibilities. Fig. 1 compares the multitask method to other approaches by visualizing mappings between sets of molecules and the quantum chemistry methods which are used to generate viable training data sets. The highest level of theory used to construct these training sets will be called the primary level. In our case, the primary level will be CCSD(T). All other levels of theory will be categorized as secondary levels. Our secondary levels will include DFT methods using different density functional approximations (DFAs). Different levels of theory may contribute data points corresponding to the same set of molecules, different sets of molecules, or some union of shared and unique sets of molecules. Furthermore, the testing data set contains primary level predictions for target molecules. Thus, the multitask method has the option to use secondary predictions for the target molecules as part of the training set. The diagrams in Fig. 1 show several options for constructing a training set for inference. The impact of these choices on inference is under-explored in literature, and the multitask method is a useful tool for investigating the benefits of different training set configurations. Crucially, the multitask framework enables training with a “dataset of opportunity,” amalgamated from existing data sets, as an alternative to expending computational time to generate entirely new data.
@@
Figure 1: Diagram of training data sets compatible with the methods considered in this work. Here, we define three sets of molecules—Core, Additional, and Target systems—and two quantum chemistry methods—primary and secondary. Our goal is to train a model that will predict our quantity of interest for the Target molecules with the accuracy of the primary method. By definition, the Core molecules are all systems for which we can compute training data with the primary method. Additional molecules are systems for which we have no primary data but which are not a part of the Target set. Theoretically, we may train using secondary predictions for molecules of all sets though not all inference methods support use of all data sets. The diagram shows the combinations of data sets and quantum chemistry methods supported by each inference method.
@@ II Statistical Models
We review the formulation of Gaussian process regression models. Subsections II.2 through II.4 discuss variants of GP regression which leverage heterogeneous sets. See references Rasmussen and Williams (2006); Kennedy and O’Hagan (2000); Leen, Peltonen, and Kaski (2012); Bonilla, Chai, and Williams (2008); Abrahamsen (1997) for more detail.
@@ II.1 Gaussian Process Regression
Our goal is to predict values of some quantity of interest (QoI) $f$, given an identifying input feature $\bm{X}$. We train our predictive model using $\{\bm{X}_{i},\bm{Y}_{i}\}_{i=1}^{n}$ where each $\bm{Y}_{i}$ is a noisy observation of $f(\bm{X}_{i})$. In our setting, we interpret imperfect first-principle predictions of the QoI as noisy observations of the true quantity. Each $\bm{X}_{i}$ is some representation of the electronic structure of a molecular system. For our numerical results, we build $\bm{X}_{i}$ using the well established Smooth Overlap of Atomic Potentials (SOAP). Bartók et al. (2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) $\bm{Y}_{i}$ is the CCSD(T) or DFT prediction of the quantity of interest for the $i^{th}$ molecular configuration.
@@
$\bm{Y}_{i}$ $=$ $\ f(\bm{X}_{i})\ +\ \varepsilon_{i},$ (1)
$\varepsilon_{i}$ $\sim$ $\ N\left(0,\ \sigma^{2}\right).$
Before inference, we choose a mean function, $\mu(\bm{X})$ to describe the expected value of $f(\bm{X})$ and a kernel function, $k(\cdot,\cdot)$, to encapsulate our assumptions about the relationship between $f(X_{i})$ and $f(X_{j})$. Typically, $k(\cdot,\cdot)$ will have hyperparameters that are selected through optimization, then fixed during inference.
@@
As a consequence of this prior assumption, any finite collection, $\{f(\bm{X}_{1}),\dots,f(\bm{X}_{n}\})$, will follow a multivariate normal distribution with mean $\bm{\mu}=\{\mu(\bm{X}_{1}),\dots,\mu(\bm{X}_{n})\}$ and covariance matrix $K$ with entries at row $i$ and column $j$ given by $k(\bm{X}_{i},\bm{X}_{j})$. We are interested in inferring the values the function takes on at target inputs, and we collect these values in the vector, $\bm{f}_{*}$. Using this notation, we write the joint distribution on the observations and targets for inference:
@@
where $I$ is the identity matrix. The specific subscripts of each covariance matrix represent the features that the matrix compares: $p$ corresponds to our primary training data and *** indicates the target set. We note that in this single-fidelity setting all training data is contained in the primary set, but this will no longer be the case when we consider surrogates informed by heterogeneous data.
@@ II.2 $\Delta$-Learning
Suppose that for each training input, we have data on our QoI from multiple sources. We thus seek an inference model that leverages $\left\{\bm{Y}_{ij}\right\}_{i=1,j=1}^{n,m}$, where we have data for molecular configurations $i=1,\dots,n$ and methods $j=1,\dots,m$. A simple “pooling” approach may put all observations into one vector and proceed with inference via Eq. (10). Then, $\bm{Y}_{ij_{1}}$ and $\bm{Y}_{ij_{2}}$ are interpreted as independent observations of the QoI, $f(\bm{X}_{i})$, for molecular configuration $\bm{X}_{i}$. The distinct behavior of each quantum chemical method is treated as a perturbation of some mean value by a realization of the noise distribution. We seek instead a model where the distribution of each $\bm{Y}_{ij}$ is sensitive to $j$. Ideally, we want to map differences in electronic structures methods to differences in their predictions.
@@
Method-specific features could be a solution, but construction of such features is challenging. One approach involves appending a vector which indicates the method category to each feature. This strategy puts a large burden on covariance hyperparameter optimization since one set of hyperparameters must encode differences between molecular configurations and between quantum chemical methods. Alternatively, to characterize the qualitative difference between methods within the features, we would have to solve the very problem for which we are creating the features in the first place—for instance, we would need to reliably predict CCSD(T) performance given CCSD(T) and DFT data from training molecules as well as DFT data for target molecules. Instead, we use this subsection to discuss $\Delta$-learning, which directly models the differences between the predictions of quantum chemistry methods.

### v2

@@ Abstract
Data generation remains a bottleneck in training surrogate models to predict molecular properties. We demonstrate that multitask Gaussian process regression overcomes this limitation by leveraging both expensive and cheap data sources. In particular, we consider training sets constructed from coupled-cluster (CC) and density functional theory (DFT) data. We report that multitask surrogates can predict at CC-level accuracy with a reduction to data generation cost by over an order of magnitude. Of note, our approach allows the training set to include DFT data generated by a heterogeneous mix of exchange-correlation functionals without imposing any artificial hierarchy on functional accuracy. More generally, the multitask framework can accommodate a wider range of training set structures—including full disparity between the different levels of fidelity—than existing kernel approaches based on $\Delta$-learning, though we show that the accuracy of the two approaches can be similar. Consequently, multitask regression can be a tool for reducing data generation costs even further by opportunistically exploiting existing data sources.
@@ I Introduction
model a data-based surrogate for quantum chemical prediction
training data possibly heterogeneous data used to construct a model
testing data data used to evaluate prediction error of a model; disjoint from training data
level of theory the first principles method used to generate quantum chemical data (e.g., CCSD(T), DFT with PBE, …)
high fidelity an accurate, expensive method such as CCSD(T)
low fidelity a method, such as DFT, which is cheaper but less accurate than the high fidelity method
task a model component trained with homogeneous data
primary the task associated with the main prediction objective of the model
secondary supplemental tasks which are included in a model to support prediction of the primary task
core (C) molecular configurations for which training data is available for the primary task
additional (A) molecular configurations for which training data is only available for secondary tasks
target (T) molecular configurations for which the model will make predictions (i.e., the testing set)
Table 1: Definitions of key terms.
@@
Predicting molecular properties based on first-principle approaches drives innovation across the physical sciences. However, exact prediction is infeasible due to the exponential scaling of the Schrödinger equation—the underlying quantum-mechanical description of materials. Decades of research have yielded numerous first-principle methods to describe electronic structures approximately, each representing a compromise between computational cost and prediction accuracy. Coupled cluster (CC) methods are typical examples of high fidelity models which offer accuracy in exchange for considerable computational cost. Specifically, CCSD(T)—the “gold standard” of quantum-chemical prediction—suffers from a steep $O(N_{e}^{7})$ scaling where $N_{e}$ is the number of electrons in the target system. Harding et al. (2008) In contrast, density functional theory (DFT) methods are popular baseline models since they provide adequate accuracy for most systems while enjoying substantially lower cost (scaling only as $O(N_{e}^{3})$). Ying, Yu, and Ying (2019); Perdew and Schmidt (2001); Goerigkab and Grimme (2011); Teale et al. (2022)
@@
Traditional workflows force users to settle on a single quantum chemical method that is cheap enough to result in manageable cost but still sufficiently high fidelity to give meaningful predictions. However, the advent of modern data-driven modeling has introduced new tools which make it possible to decouple the step of estimating the properties of unseen molecular systems from the step of running expensive first-principles computations. Statistical surrogate models are trained with predictions from first-principle methods and, once trained, can be substituted into workflows, making the prediction step considerably cheaper. In the context of building potential energy surfaces and interatomic potentials, these data-driven approaches have become the de facto standard. Smith et al. (2019); Dral et al. (2020); Goodlett, Turney, and Schaefer (2023); Zaverkin et al. (2023) Simultaneously, large-scale data generation is facilitated by sophisticated workflows for high-throughput calculations Curtarolo et al. (2012); Jain et al. (2011); Huber et al. (2020) and robust DFT methods inspired by mathematical research. Teale et al. (2022); Cancès et al. (2022); Herbst and Levitt (2020, 2022); Cancès et al. (2023) Increasing amounts of data are openly available Ruddigkeit et al. (2012); Ramakrishnan et al. (2014); Chanussot* et al. (2021); Smith, Isayev, and Roitberg (2017) and are, in principle, ready for incorporation into statistical surrogates, but quality is generally not homogeneous as each dataset’s author has made different methodological choices. Consequently, most surrogate models are single fidelity: all training data is generated using the same quantum-chemical method and the same numerical approach. Heterogeneity has thus been an obstacle to opportunistically exploiting the already available data. In this work, we construct surrogates which do not require a single choice between accurate (but expensive) and fast (but crude) methods, but instead combine the advantages of both. Specifically, we consider multitask methods which construct a relationship between multiple regression problems without requiring precise information about the respective accuracy of each source. Bonilla, Chai, and Williams (2008); Leen, Peltonen, and Kaski (2012)
@@
Figure 1: Diagram of training data sets compatible with the methods considered in this work. Here, we define three sets of molecules—core, additional, and target systems—and two fidelities—primary and secondary. Our goal is to train a model that will predict our quantity of interest for the target molecules with the accuracy of the primary method. The molecular space covered by secondary training data is a design choice for which we present six options, labeled C, CT, A, AT, CA, and CAT. Key terms are defined in Table 1.
Our multitask framework can accommodate a range of heterogeneous training set structures. For example, suppose we aim to design a surrogate which predicts the three-body energy for any configuration of a water trimer with accuracy comparable to CCSD(T). We may only have access to $100$ CCSD(T) data points, all corresponding to configurations with HOH angles smaller $110$ degrees. Additional data may be provided by including a more exhaustive dataset based on DFT simulations with the PBE functional. Notably, even though the geometries of the molecular configurations in both datasets may not match, they can both be fully employed to train a multitask surrogate. Moreover, if higher predictive accuracy is desired, further training data can be created using yet another quantum-chemical method. For example, we may employ BLYP-based calculations on all molecular configurations covered by CCSD(T) and PBE, plus some configurations which we know we would eventually target for prediction. Remarkably, in this case, all available data can be used without imposing any assumption about which of the DFT methods (here, BLYP or PBE) is more accurate for the considered systems. Internally, a multitask surrogate creates a statistical submodel for each of the categories of data we supply it with (here, different levels of theory, i.e., CCSD(T), PBE, BLYP). In the language used in the rest of the article we call the submodel associated to our main prediction objective—the three-body energy at CCSD(T) level—the primary task. The other submodels—here associated to the DFT levels of theory— strive to improve the predictive accuracy of the primary task and are called secondary tasks. For a concise overview of these and other terms employed in the article, see Table 1.
In this work, we will investigate such multitask models and the impact of training set design (i.e., the distribution of training data between the primary and secondary tasks) on prediction accuracy. Fig. 1 visualizes the different training set structures which we consider. Secondary task training sets may share some molecular configurations with primary task training sets, include configurations unique to the secondary task, and may even include configurations from the testing data set. An example of the latter case is the use of a BLYP energy prediction at the same geometry where we later apply the surrogate to predict the CCSD(T) energy. The impact of these choices on inference is under-explored in literature, and multitask models are useful tools for investigating the benefits of different training set configurations. Crucially, the multitask framework enables training with a “dataset of opportunity,” amalgamated from existing data sets, as an alternative to expending computational time to generate entirely new data.
In this paper, we construct submodels for each task using Gaussian process (GP) regression, a popular tool for constructing surrogates informed by heterogeneous data. Several works Pilania, Gubernatis, and Lookman (2017); Batra et al. (2019); Patra et al. (2020) have explored the use of multifidelity GP models Kennedy and O’Hagan (2000) trained using two fidelities, namely a high and a low fidelity DFT approach. Alternatively, $\Delta$-learning may be employed to model the difference between two levels of theory, then by adding that difference to low level predictions of the target to produce high level predictions. Ramakrishnan et al. (2015); Dral, Zubatiuk, and Xue (2023); Batra et al. (2019) Though GP regression and the related technique of kernel ridge regression are common choices for fitting such models, in principle, any method of fitting a surrogate may be employed. Batra et al. (2019); Dral, Zubatiuk, and Xue (2023). Hierarchical machine learning (hML) models extend the $\Delta$-learning framework to leverage more than two training data sets by learning and summing multiple difference models. Dral et al. (2020); Dral, Zubatiuk, and Xue (2023) However, between each pair of training sets, data points must be in some way aligned. For example, in a model of the difference between DFT and CCSD(T), DFT predictions may only contribute to the training set if the corresponding CCSD(T) prediction is also available. Further, when either hML or multifidelity GP regression learn from more than two heterogeneous training sets, these sets must be assigned some order—for example, following a hierarchy in accuracy. Recent work Vinod, Kleinekathöfer, and Zaspel (2023); Vinod et al. (2023) has proposed an optimized linear combination of submodels as an alternative, but all coefficients in this linear combination must be re-optimized whenever any submodel is altered, and opportunistic data set construction is not fully explored.
Progress towards multifidelity approaches has also been made in the context of neural network surrogates. For example, transfer learning methods train an initial neural network on a larger secondary dataset, then re-optimise part of the parameters using a primary dataset. Towell and Shavlik (1994); Fu (1989); Smith et al. (2019) For the prediction of atomistic properties, the secondary dataset might contain millions of DFT calculations while the primary might be a smaller set of CCSD(T) predictions. In prior studies, it has been observed that transfer learning can produce higher accuracy than training a network on only one of these datasets. Dral et al. (2020); Dral, Zubatiuk, and Xue (2023) While the framework we discuss in this work is not restricted to GP regression strategies, we will focus exclusively on GP methods. In comparison to neural network approaches, where a rigorous mathematical explanation for the low generalization error of well-trained models is still lacking Bartlett, Montanari, and Rakhlin (2021); Lotfi et al. (2022), GP approaches are more amenable to mathematical analysis. Rasmussen and Williams (2006) This comparative transparency simplifies interpretation of the surrogate’s predictive process in the diverse training data configurations we analyze. Additionally, GP methods have the advantage of yielding accurate predictions even for relatively small training set sizes. While the training cost of GP regression may become large as the size of the training set increases, many strategies for efficient scaling have been developed and can be routinely employed.Quinonero-Candela and Rasmussen (2005); Liu et al. (2020); Wilson, Dann, and Nickisch (2015); Cole, Christianson, and Gramacy (2021)
@@ II Statistical Models
We review the formulation of Gaussian process regression models. Subsections II.2 through II.5 discuss variants of GP regression which leverage heterogeneous data sets. See references Rasmussen and Williams (2006); Kennedy and O’Hagan (2000); Leen, Peltonen, and Kaski (2012); Bonilla, Chai, and Williams (2008); Abrahamsen (1997) for more detail.
@@ II.1 Gaussian Process Regression
Our goal is to predict values of some quantity of interest (QoI) $f$, given an input feature vector $\bm{X}_{m}$ describing the $m^{th}$ molecular configuration. We train our predictive model using $\{(\bm{X}_{m},\bm{Y}_{m})\}_{{m}=1}^{M}$ where each $\bm{Y}_{m}$ is a noisy observation of $f(\bm{X}_{m})$. In our setting, we interpret imperfect first-principle predictions of the QoI as noisy observations of the true quantity. Each $\bm{X}_{m}$ is some representation of the electronic structure of a molecular system. For our numerical results, we build $\bm{X}_{m}$ using the well established Smooth Overlap of Atomic Potentials (SOAP). Bartók et al. (2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) $\bm{Y}_{m}$ is the CCSD(T) or DFT prediction of the quantity of interest for the ${m}^{th}$ molecular configuration.
@@
$\bm{Y}_{m}$ $=$ $\ f(\bm{X}_{m})\ +\ \varepsilon_{m},\quad m=1,\dots,M,$ (1)
$\varepsilon_{m}$ $\sim$ $\ N\left(0,\ \sigma^{2}\right).$
Before inference, we choose a mean function $\mu(\bm{X}_{m})$ to describe the expected value of $f(\bm{X}_{m})$ and a kernel function, $k(\cdot,\cdot)$, to encapsulate our assumptions about the relationship between $f(\bm{X}_{m})$ and $f(\bm{X}_{m^{\prime}})$. Typically, $k(\cdot,\cdot)$ will have hyperparameters that are selected through optimization, then fixed during inference.
@@
As a consequence of this prior assumption, any finite collection, $\{f(\bm{X}_{1}),\dots,f(\bm{X}_{M})\}$, will follow a multivariate normal distribution with mean $\bm{\mu}=\{\mu(\bm{X}_{1}),\dots,\mu(\bm{X}_{M})\}$ and covariance matrix $K$ with entries at row $m$ and column $m^{\prime}$ given by $k(\bm{X}_{m},\bm{X}_{m^{\prime}})$. It is common practice to center the observation vector and set $\bm{\mu}=\bm{0}$. We retain a more general formulation here to emphasize that it is possible to construct a feature-dependent prior mean.
We are interested in inferring the values the QoI takes on at target inputs, and we collect these values in the vector, $\bm{f}_{*}$. Using this notation, we write the joint distribution of the observations and targets for inference:
@@
where $I$ is the identity matrix. The specific subscripts of each covariance matrix represent the features that the matrix compares: $p$ corresponds to our primary training data and * indicates the target set. We note that in this single task setting, all training data is contained in the primary set, but this will no longer be the case when we consider surrogates informed by heterogeneous data.
@@ II.2 Multiple Observation Sets
Suppose we have access to heterogeneous training data which includes predictions of our QoI by multiple first principles methods. We thus seek an inference model that leverages all data $\bigcup_{n=1}^{N}\{\bm{Y}_{mn}\}_{m\in\mathcal{M}_{n}}$, where each $n=1,\dots,N$ represents a task and the corresponding set $\mathcal{M}_{n}$ contains indices of the molecular configurations available to train a model for task $n$. A simple “pooling” approach could put all observations into one vector and proceed with inference via Eq. (4). Then, if $m\in\mathcal{M}_{n}\cap\mathcal{M}_{n^{\prime}}$, we can interpret $\bm{Y}_{mn}$ and $\bm{Y}_{mn^{\prime}}$ as independent observations of the same QoI, $f(\bm{X}_{m})$, for molecular configuration $\bm{X}_{m}$. The distinct behavior of each quantum chemical method is treated as a perturbation of some mean value by a realization of the noise.
@@
Rather than doing this, we seek a model where the distribution of each $\bm{Y}_{mn}$ is sensitive to $n$. Ideally, we want to map differences in electronic structure methods to differences in their predictions. Method-specific features could be a solution, but construction of such features is challenging. One approach involves appending a vector which indicates the method category to each feature. This strategy puts a large burden on covariance hyperparameter optimization since one set of hyperparameters must encode differences between molecular configurations and between quantum chemical methods. Alternatively, to characterize the qualitative difference between methods within the features, we would have to solve the very problem for which we are creating the features in the first place: for instance, we would need to reliably predict CCSD(T) performance given CCSD(T) and DFT data from training molecules as well as DFT data for target molecules.
Instead, we will extend the framework of GP regression to model relationships between multiple prediction tasks.
II.3 $\Delta$-Learning